The two plates of a capacitor hold +2500 μC and -2500 μC of charge, respectively, when the potential difference is 940 V .

What is the capacitance?
Answer #1

Solution)

Given,

Charge, Q=2500*10^-6 C

Voltage, V=940V

We know, Capacitance, C=Q/V=(2500*10^-6)/(940)

C=2.65*10^-6 F (Ans)

===========